# Sigma: Scalable Spectral Insights for LLM Collapse

### v1

@@ C.4 Surface-form proxy diagnostics and alignment with Sigma-UB
Endpoint comparison and co-movement.
To quantify the qualitative overlay, Table 14 reports proxy endpoints ($g=0$ vs. $g=50$) and simple correlation diagnostics between Track II drift and the proxies over generations. The key operational takeaway is that true recursion (S2) yields substantially larger surface-form concentration (HHI increases by nearly three orders of magnitude relative to its $g=0$ baseline) than restart-from-base (S1), and this increase co-varies strongly with Track II under a rank-based measure.
@@
Table 14: Surface-proxy alignment with Track II (FIN). Distinct-2 decreases and hashed $n$-gram HHI increases in both settings, with substantially stronger concentration under true recursion (S2). Correlations are computed over generations using the per-generation series (including $g=0$). In S1, $\Delta U$ and distinct-2 co-move strongly (both decrease), yielding a high positive Pearson correlation. In S2, distinct-2 is noisier (and can transiently increase despite strong contraction), while HHI retains a strong monotone association with $\Delta U$ (Spearman $\approx-0.69$), consistent with the overlay in Figure 3.
Full per-generation proxy table (optional).
For completeness and auditability, the full per-generation proxy values can be included as a table. This is often useful when debugging changes to tokenization, hashing, or evaluation decoding.
C.5 Within-domain localization by frozen prompt buckets (optional)
Because the evaluation prompt bank is labeled by prompt type (creative, divergent, analogy/ELI5, what-if, neutral), we can localize contraction to specific regimes while keeping the evaluation distribution fixed over time. Concretely, we re-run the evaluation pipeline but restrict to prompts in a single bucket, generating one response per prompt under the same decoding, embedding with the same frozen encoder, and computing the same Sigma-UB drifts. This is a slice of a fixed evaluation set, not a moving target, so it does not introduce evaluation drift.
Figure 4: Bucket localization under S1 (FIN; optional). Sigma-UB drifts computed after restricting evaluation to each frozen prompt bucket. Bucket curves can be noisier than the aggregate because each bucket contains fewer prompts; we therefore interpret bucket localization as a qualitative “where does collapse concentrate?” tool rather than as a primary metric.
Table 15: Track-II drift by bucket under S1 (FIN; optional). Each entry is $\Delta U_{\mathrm{LLN,cov}}(\delta)$ computed using only prompts from the given bucket (baseline $g=0$). In this run, Creative tends to exhibit the largest-magnitude negative drift by late generations, while other buckets are comparatively less contracted. Because bucket sizes differ, cross-bucket comparisons should be interpreted qualitatively unless normalized for prompt count.
@@ C.5 Within-domain localization by frozen prompt buckets (optional)
Bucket localization helps answer a practical question: is contraction uniformly distributed across prompt types, or concentrated where the model is pushed into higher-entropy regimes? In the reported S1 run, the Creative bucket shows consistently larger-magnitude Track-II contraction than the aggregate, suggesting that open-ended prompts amplify the representational narrowing induced by synthetic recursion. That said, bucket series are noisier (and can be non-monotone) because each bucket contains fewer prompts and fewer embedded blocks, so we use them as a diagnostic slice rather than as a headline number.

### v2

@@ C.5 Surface-form proxy diagnostics and alignment with Sigma-UB
To quantify the qualitative overlay, Table 11 reports proxy endpoints ($g=0$ vs. $g=50$) and simple correlation diagnostics between Track II drift and the proxies over generations. The key operational takeaway is that true recursion (S2) yields substantially larger surface-form concentration than restart-from-base (S1), and this increase co-varies strongly with Track II under a rank-based measure.
@@
Table 11: Surface-proxy alignment with Track II (TECH). Distinct-2 decreases and hashed $n$-gram HHI increases in both settings, with substantially stronger concentration under true recursion (S2). Correlations are computed over generations using the per-generation series (including $g=0$). In S1, $\Delta U$ and distinct-2 co-move strongly (both decrease), yielding a high positive Pearson correlation. In S2, distinct-2 is noisier (and can transiently increase despite strong contraction), while HHI retains a strong monotone association with $\Delta U$ (Spearman $\approx-0.69$), consistent with the overlay in Figure 3.
Figure 4: Bucket localization under S1 (TECH; optional). Sigma-UB drifts computed after restricting evaluation to each frozen prompt bucket. Bucket curves can be noisier than the aggregate because each bucket contains fewer prompts; we therefore interpret bucket localization as a qualitative “where does collapse concentrate?” tool rather than as a primary metric.
C.6 Within-domain localization by frozen prompt buckets
Because the evaluation prompt bank is annotated by prompt type (creative, divergent, analogy/ELI5, what-if, neutral), we can attribute model collapse to particular regimes while holding the evaluation framework constant over time. Specifically, we modularize the evaluation pipeline by limiting the framework to prompts from a single category, producing one response per prompt with identical decoding settings, embedding these responses using the same frozen encoder, and then computing the same Sigma-UB drift metrics.
@@ C.6 Within-domain localization by frozen prompt buckets
Bucket localization helps answer a practical question: is model collapse uniformly distributed across prompt types, or concentrated where the model is pushed into higher-entropy regimes? In the reported S1 run, the Creative bucket shows consistently larger-magnitude Track-II contraction than the aggregate, suggesting that open-ended prompts amplify the representational collapse induced by synthetic recursion. That said, bucket series are noisier because each bucket contains fewer prompts and therefore fewer sample size.
Table 12: Track-II drift by bucket under S1 (TECH; optional). Each entry is $\Delta U_{\mathrm{LLN,cov}}(\delta)$ computed using only prompts from the given bucket (baseline $g=0$). In this run, Creative tends to exhibit the largest-magnitude negative drift by late generations, while other buckets are comparatively less contracted. Because bucket sizes differ, cross-bucket comparisons should be interpreted qualitatively unless normalized for prompt count.